# Towards Better Comprehension of Breaking Changes in the NPM Ecosystem

### v1

@@ 4.3. RQ3: How do Developers Make Behavioral Breaking Change?
  1. (1)
    Changing the specification of return values (79/1,034 cases). One typical case is that the type of returned objects are changed. For example, in commit b99f6d3 of automattic/mongoose, MongooseArray.map() returned a plain JavaScript array rather than a headless Mongoose array:
@@
    - const ret = super.map.apply(this, arguments);
    - ret[arraySchemaSymbol] = null;
    - ret[arrayPathSymbol] = null;
    - ret[arrayParentSymbol] = null;
@@
  2. (2)
    Changing the process of some options (231/1,034 cases). Suppose that one possible value of an option (can be in JSON configuration or parameters) is not supported, then the code that gets properties from the option will be changed, and the corresponding branch of a specific option value will be removed. For example, in commit ad4f1493 of project octokit/rest.js [^18] https://github.com/octokit/rest.js/commit/ad4f1493 (shown below), the judgment of options.type is changed, and netrc branch in a switch structure is also removed to completely remove support for netrc mechanism.
@@
    - ’basic|oauth|client|token|integration|netrc’.indexOf(options.type) $=$ -1
@@
    - ’basic’, ’integration’, ’oauth’, ’client’ or ’netrc’"
    + ’basic’, ’integration’, ’oauth’, or ’client’"
    In CLI tool projects, developers often change the process of some options to remove or change the purpose of an option. For example, in commit f6fd0c3 of pnpm/pnpm, the option --store-path is not an alias of --store anymore. Therefore, developers remove the code of handling the option:
@@
    - if (opts.storePath && !opts.store) {
    - logger.warn(’the ‘store-path‘ config is deprecated. Use ‘store‘ instead.’)
    - opts.store = opts.storePath
@@
  3. (3)
    Changing the default behavior for unprovided values (203/1,034 cases). Typically, in commit 72bbda7f of npm/cli [^19] https://github.com/npm/cli/commit/72bbda7f (the diff shown below), the default value of local variable “depthToPrint” is set to zero while it is initially “undefined”.
    const { /* ... */, depth, /* ... */ } = npm.flatOptions;
    const depthToPrint = all ? Infinity : (depth || 0);
@@
  4. (4)
    Changing error handling process (42/1,034 cases). One case is changing change the content or format of the error objects. For example, in commit dd6306e3 of octokit/rest.js [^20] https://github.com/octokit/rest.js/commit/dd6306e3, the error message is parsed as a JSON object and the properties will be put into error object (the existing message property will be overwritten). Since developers usually parse error.message, this change will make the code like JSON.parse(error.message) not work.
@@
    Another case is changing the criteria of detecting errors, e.g, some status values are no longer regarded as success. In the commit 201f189d in angular/angular [^21] https://github.com/angular/angular/commit/201f189d, if the return code of an XHR (XML HTTP Request) is not 200, an error will occur, while before the code change, if the return codes between 200 and 300 (exclusive) were all regarded as success states.
    // Breaking Change in xhr_backend.ts
@@
    // in http_util.ts
@@
We also compare the behavioral BCs those in Maven ecosystem. Zhang et al. have studied the behavioral BC types in Maven projects (Zhang et al., 2022). They found that changing execution logic and changing calculation of the output make up over 60% of the total BCs collected by them. They also studied benign changes that will not cause BCs, such as additional/changed/deleted conditions and branches (over 55% of the total benign changes), additional try/catch statements, and assignment revision. However, in our findings about source code patterns of JavaScript behavioral BCs, it is very common to change conditions in “if” statements to remove support for some option values or to change evaluation criteria of some status, although it seems benign. Hence, in JavaScript, the condition changes that involve option/configuration values deserve attention since they are more probable to be BCs.
@@ 4.4. RQ4: Why do Developers Make Breaking Changes in NPM Ecosystem?
In this section, we summarize six major motivations behind making BCs by carefully analyzing our collected BCs. We show the connection between BC types and motivations in Figure 3.
@@
  • To unify the approach of an functionality (21 cases). The functionality can be achieved by other methods, or by using other packages. Therefore, the current implementation can be removed. For example, in commit 1354171 of thi-ng/umbrella, developers removed function “writeFile”, and recommend using “writeFile” provided by another package @thi-ng/rstream-log-file since the third-party function can also achieve the same functionality.
  • To deprecate and remove an outdated functionality (413 cases). Some outdated input (e.g., authentication mechanism) are deprecated thus developers remove their support. For example, in commit ad4f1493 of Rest.js, the support for netrc authentication method was removed while others were retained.
  • To better organize the project (35 cases). Moving part of the functionality out and create a new package will make the project better organized. For example, in commit 0fa2830222222https://github.com/jsdoc/jsdoc/commit/0fa2830 of jsdoc/jsdoc, the package @jsdoc/core was reorganized and some methods in @jsdoc/core moved to a new package @jsdoc/cli.
@@
  • To make the identifier name reflect the actual functionality (68 cases). The name of a class, interface or method, etc., should reflect the actual functionality. Therefore, after a code change, the identifier names tend to be modified. For example, in issue 3403 of angular/angular [^23] https://github.com/angular/angular/issue/3403, the developer complaint that “When I first saw the requestCheck() method on ChangeDetectorRef, I assumed it was how I manually run change detection. Others have assumed this as well”. Hence after the issue was reported, the identifier requestCheck is changed to markForCheck.
@@
  • To fix typos (2 cases). Simply fixing a typo. For example, renaming “MetricAarmProps” to “MetricAlarmProps”.
@@ 5. Implications
Automatic naming and renaming techniques can be applied in NPM projects. In Section 4.3, we conclude that the main reason for renaming identifiers is to make identifier name reflect the actual functionality. Therefore, on the one hand, poor identifier name is a noticeable technical debt in JavaScript projects, hence automatic naming can be utilized to provide good identifier names, and on the other hand, in some cases, renaming identifiers is reasonable due to the necessary adaptation to frequent code changes in NPM ecosystem, thus we suggest applying automatic renaming techniques to make identifier names adapt to code evolution.
@@
  • They are not code-change-aware. Since we have found that many a proportion of Rename BCs are performed to ensure code consistency (Figure 3 in Section 4.3), a renaming technique had better leverage knowledge in prior related code changes to recommend proper names.
Tools for detecting code with similar functionalities are necessary. According to Figure 3, a number of Remove BCs are motivated by approach for unification, i.e., provide a unified API for a specific functionality. However, unlike the casualty changes (e.g., removing a class then the code referring to it should also be removed), the relationship between these co-occurring changes is indirect and not simple to detect. Besides, in our BC extraction process, we find developers make other small changes for better performance, compatibility, and robustness though they are unrelated to BCs. This also to some degree interferes with BC identification. The findings in our investigation can provide the following directions for future research:
@@
  • Distinguish non-breaking changes from breaking changes and casualty changes to provide more convenience for developers.
Future research can focus on detecting more types of behavioral BCs. Syntactic BC detection (e.g., Rename, Remove and Change Signature) is somehow trivial and many refactoring detection-based tools can achieve it. For example, APIDiff (Brito et al., 2018), Clirr (Cli, [n. d.]), Japicmp (jap, [n. d.]) and RevAPI (Rev, [n. d.]) can be used for Java projects, PyCompat (Zhang et al., 2020) and Pidiff (pid, [n. d.]) can be applied in for Python libraries, and a variant of APIDiff can be used for JavaScript to characterize refactoring history (Brito et al., 2021). Hence the future research works should focus on behavioral BCs that cannot be simply identified. Existing such approaches for JavaScript are NoRegrets (Mezzetti et al., 2018) and its enhanced version NoRegrets+ (Møller and Torp, 2019). In Section 1, we have highlighted their reliance on test cases in downstream projects. In addition, the two approaches can only detect type-related BCs. Specifically, in the first phase, they run client test cases and monitor the flows of values by program instrumentation to build API models (intuitively, type restriction of APIs), and in the second phase, they rerun the client test cases with updated upstream code to check whether type restrictions are changed during execution. Therefore, they cannot detect non-type-related BCs, e.g., a BC that changes the semantics of the returned string, while the return type is still string. According to Section 4.3, there are five types of behavioral BC actions, where only changing the specifications of return values can directly cause type changes, while other behavioral change may not change object types during execution. Therefore, we recommend future works try to deal with various types of behavioral BCs uncovered by this study.
@@ 6. Threats to Validity
Threats to internal validity are mainly related to the clarity of developers’ documentation of breaking changes. Few developers might have not explicitly documented the BCs in commit messages, and commit messages (as well as text in issues and pull requests) might not fully reflect developers’ reason for making BCs. Also, the discussion and documentation related to BCs may not fully reflect developers’ opinions. Second, in our manual analysis, we might still not understand developers’ intentions since many BC commit messages do not clearly explain what and why they make BCs. Additionally, we might regard some changes in code diff as BC actions since they are confusing. To address this potential threat, we have implemented a double-checking process to verify the accuracy of our analysis results.
Threats to construct validity are concerned with the errors during breaking change selection. While checking the compliance to Conventional Commits standard, we simply use the regular expression, which might incorrectly classify some commit messages. Also, in the process of selecting commits for further study (Section 3.3), we might abandon some commits, although they satisfy our selection criteria.
Threats to external validity refer to the generalizability of our study. In our study, while we involve 1,519 BCs extracted from 381 randomly sampled projects from a vast number of popular NPM projects, and these projects cover diverse application fields of JavaScript language, such as utilities, frontend, Web APIs, database, etc., the extracted BCs only account for a small proportion of BCs in the whole NPM ecosystem. Besides, in this study we do not consider less popular projects since they have short development history and it is less probable to yield some patterns from these projects. However, the BC practices in less popular projects may still differ from those in popular projects since they are not so well maintained.
Threats to conclusion validity relate to insufficient high-quality BC commits to support our findings. To mitigate the threats, we try our best to extract breaking changes from commit messages, issues, pull requests and changelogs to construct our BC dataset. We also carefully inspect our collected BCs and retain those related to JavaScript source code changes and explicitly documented by developers, which can help increase the quality of our dataset.
@@ 7.3. Breaking Change Detection and Analysis on Other Platforms
Besides JavaScript, a lot of studies focus on detecting and analyzing BCs for other programming languages, especially Java and Python, and their techniques may be learned by BC detection for JavaScript. Brito et al. presented APIDiff (Brito et al., 2018) that can detect syntax-related breaking changes in Maven projects, such as method removal and visibility loss by reusing the refactoring detection tool RefDiff (Silva and Valente, 2017). Some open-source tools can also check Java syntactic breaking changes, such as Clirr (Cli, [n. d.]) and RevAPI (Rev, [n. d.]). For Python language, Du et al. proposed AexPy (Du and Ma, 2022) that can detect similar types of breaking changes like module removal and addition of required parameters, which extends the existing tool PyCompat (Zhang et al., 2020) and Pidiff (pid, [n. d.]). Regarding non-syntactic breaking changes, to the best of our knowledge, Zhang et al. proposed Sembid (Zhang et al., 2022) to detect behavioral BCs by measuring the semantic difference of call graphs between old and new programs: if a code change’s semantic difference is larger than a threshold and not identified as benign change, it is classified as a BC. However, their used semantic diffs reflect the structural changes of programs, while many semantic BCs do not need many of those changes, e.g., only changing one condition, and adding a small change to the output string. Additionally, as the authors have pointed out, Sembid cannot detect high-level BCs such as re-implementation of a method. Therefore, directly adapting Sembid to JavaScript projects is not suitable.
Additionally, with the help of BC detection techniques, many works analyzed the impacts of breaking changes to downstream projects. For example, Jayasuriya et al. (Jayasuriya et al., 2023, 2024) investigated the impacts of breaking changes (especially behavioral breaking changes) to client applications in Maven ecosystem. They concluded with many findings, e.g., 11.58% of the dependency updates contain breaking changes that impact the clients and 2.30% version updates have behavioral breaking changes that impacted client tests. Similar to the findings in Venturini et al.’s study (Venturini et al., 2023), while the breaking changes are only detected in a small number of code changes or version updates, they still cause negative impacts to downstream projects.

### v2

@@ 4.3. RQ3: How do Developers Make Behavioral Breaking Change?
[…]
Changing the specification of return values (79 cases). One typical case is that the type of returned objects are changed. For example, in commit b99f6d3 of automattic/mongoose, the method MongooseArray.map() returned a plain JavaScript array rather than a headless Mongoose array:
@@
- const ret = super.map.apply(this, arguments);
- ret[arraySchemaSymbol] = null;
- ret[arrayPathSymbol] = null;
- ret[arrayParentSymbol] = null;
@@
[…]
Changing the process of some options (231 cases). Suppose that one possible value of an option (can be in JSON configuration or parameters) is not supported, then the code that gets properties from the option will be changed, and the corresponding branch of a specific option value will be removed. For example, in commit ad4f1493 of project octokit/rest.js [^19] https://github.com/octokit/rest.js/commit/ad4f1493 (shown below), the judgment of options.type is changed, and netrc branch in a switch structure is also removed to completely remove support for netrc mechanism.
@@
- ’basic|oauth|client|token|integration|netrc’.indexOf(options.type) $=$ -1
@@
- ’basic’, ’integration’, ’oauth’, ’client’ or ’netrc’")
+ ’basic’, ’integration’, ’oauth’, or ’client’")
In CLI tool projects, developers often change the process of some options to remove or change the purpose of an option. For example, in commit f6fd0c3 of pnpm/pnpm, the option --store-path is not an alias of --store anymore. Therefore, developers remove the code of handling the option, shown as follows:
@@
- if (opts.storePath && !opts.store) {
- logger.warn(’the store-path config is deprecated.’)
- opts.store = opts.storePath
@@
// more code ...
[…]
Changing the default behavior for unprovided values (203 cases). The arguments of a function can often affect the behavior. When the value of a parameter is not provided (i.e., being undefined), developers may design special program logic to deal with this case, or make the parametre still remain undefined. Sometimes the default program logic of dealing with such case can be changed. As an example, in commit 72bbda7f of npm/cli [^20] https://github.com/npm/cli/commit/72bbda7f, the default value of local variable depthToPrint is set to zero while it is initially undefined. After this change, the code is as follows:
1const { /* ... */, depth, /* ... */ } = npm.flatOptions;
2const depthToPrint = all ? Infinity : (depth || 0);
3// more code
@@
[…]
Changing error handling process (42 cases). One case is changing change the content or format of the error objects. For example, in commit dd6306e3 of octokit/rest.js [^21] https://github.com/octokit/rest.js/commit/dd6306e3, the error message is parsed as a JSON object and the properties will be put into error object (the existing message property will be overwritten). Since developers usually parse error.message, this change will make the code like JSON.parse(error.message) not work. The code diff below shows this change:
@@
Another case of this category is changing the criteria of detecting errors, e.g, some status values are no longer regarded as success. In the commit 201f189d in angular/angular [^22] https://github.com/angular/angular/commit/201f189d, if the return code of an XHR (XML HTTP Request) is not 200, an error will occur, while before the code change, if the return codes between 200 and 300 (exclusive) were all regarded as success states. The breaking change in xhr_backend.ts is shown as follows:
@@
The code in http_util.ts that contains the new anonymous function isSuccess:
@@
We also compare the behavioral breaking changes to those in the Maven ecosystem. Zhang et al. have studied the behavioral breaking change types in Maven projects (Zhang et al., 2022). They found that changing execution logic and changing calculation of the output make up over 60% of the total breaking changes collected by them. They also studied benign changes that will not cause incompatibilities, such as additional/changed/deleted conditions and branches (over 55% of the total benign changes), additional try-catch statements, and assignment revision. However, in our findings about source code patterns of JavaScript behavioral breaking changes, it is very common to change conditions in “if” statements to remove support for some option values or to change evaluation criteria of some results, although it seems benign. Hence, in JavaScript, the condition changes that involve option/configuration values deserve attention since they are more probable to be breaking changes.
@@ 4.4. RQ4: Why do Developers Make Breaking Changes in NPM Ecosystem?
In this section, we summarize six major motivations behind making breaking changes by carefully analysis of our collected breaking changes. We show the connection between BC types and motivations in Figure 3.
@@
  • To unify the approach of a functionality (21 cases). The functionality can be achieved by other methods, or by using other packages. Therefore, the current implementation can be removed. For example, in commit 1354171 of thi-ng/umbrella, developers removed function writeFile, and recommend using writeFile provided by another package @thi-ng/rstream-log-file since the third-party function can also achieve the same functionality.
  • To deprecate and remove an outdated functionality (413 cases). Some outdated input (e.g., authentication mechanism) are deprecated thus developers remove their support. For example, in commit ad4f1493 of octokit/rest.js, the support for netrc authentication method was removed while others were retained.
  • To better organize the project (35 cases). Moving part of the functionality out and create a new package will make the project better organized. For example, in commit 0fa2830232323https://github.com/jsdoc/jsdoc/commit/0fa2830 of jsdoc/jsdoc, the package @jsdoc/core was reorganized and some methods in @jsdoc/core moved to a new package @jsdoc/cli.
@@
  • To make the identifier name reflect the actual functionality (68 cases). The name of a class, interface or method, etc., should reflect the actual functionality. Therefore, after a code change, the identifier names tend to be modified. For example, in issue #3403 of angular/angular [^24] https://github.com/angular/angular/issue/3403, the developer complaint that “When I first saw the requestCheck() method on ChangeDetectorRef, I assumed it was how I manually run change detection. Others have assumed this as well”. Hence after the issue was reported, the identifier requestCheck is changed to markForCheck.
@@
  • To fix typos (2 cases). For example, renaming MetricAarmProps to MetricAlarmProps.
@@ 5. Implications
Automatic naming and renaming techniques can be applied in NPM projects. In Section 4.3, we conclude that the main reason for renaming identifiers is to make identifier name reflect the actual functionality. Therefore, on the one hand, poor identifier naming is a noticeable technical debt in JavaScript projects, hence automatic naming can be utilized to provide good identifier names, and on the other hand, in some cases, renaming identifiers is reasonable due to the necessary adaptation to frequent code changes in NPM ecosystem, thus we suggest applying automatic renaming techniques to make identifier names adapt to code evolution.
@@
  • They are not code-change-aware. Since we have found that many a proportion of Rename breaking changes are performed to ensure code consistency (Figure 3 in Section 4.3), a renaming technique had better leverage knowledge in prior related code changes to recommend proper names.
Tools for detecting code with similar functionalities are necessary. According to Figure 3, a number of Remove breaking changes are motivated by approach for unification, i.e., provide a unified API for a specific functionality. However, the relationship between these two or more code changes is indirect and not simple to detect. Besides, in our BC extraction process, we find developers make other small changes for better performance, compatibility, and robustness though they are unrelated to breaking changes. This also to some degree interferes with BC identification. The findings in our investigation can provide the following directions for future research:
@@
  • Distinguish non-breaking changes from breaking changes and ordinary code changes to provide more convenience for developers.
Future research can focus on detecting more types of behavioral breaking changes. Existing approaches for JavaScript behavioral breaking change detection are NoRegrets (Mezzetti et al., 2018) and its enhanced version NoRegrets+ (Møller and Torp, 2019), to the best of our knowledge. In Section 1, we have highlighted their reliance on test cases in downstream projects. In addition, the two approaches can only detect type-related breaking changes. Specifically, in the first phase, they run client test cases and monitor the flows of values by program instrumentation to build API models (intuitively, type restriction of APIs), and in the second phase, they rerun the client test cases with updated upstream code to check whether type restrictions are changed during execution. Therefore, they cannot detect non-type-related breaking changes, e.g., a BC that changes the semantics of the returned string, while the return type is still string. According to Section 4.3, there are four types of behavioral BC actions, where only changing the specifications of return values can directly cause type changes, while other behavioral changes may not change object types during execution. Therefore, we recommend future works try to deal with various types of behavioral breaking changes uncovered by this study.
@@ 6. Threats to Validity
Threats to internal validity are mainly related to the clarity of developers’ documentation of breaking changes. Few developers might have not explicitly documented the breaking changes in commit messages, and commit messages (as well as text in issues and pull requests) might not fully reflect developers’ reason for making breaking changes. Also, the discussion and documentation related to breaking changes may not fully reflect developers’ consideration. Second, in our manual analysis, we might still not understand developers’ intentions since many breaking changes’ commit messages do not clearly explain what and why they make breaking changes. Additionally, we might regard some changes in code diff as BC-related actions since they are confusing. To address this potential threat, we have double-checked the analysis results to ensure the quality.
Threats to construct validity are concerned with the errors during breaking commit selection. In our study, we mainly consider the commits that follow Conventional Commits specification and collect breaking change commits that contain BREAKING CHANGE token, which may miss the commits that not fully follow the specification but are actually breaking. To mitigate this, we try our best to collect breaking changes from software documentation, especially changelogs, issues and pull requests (mentioned in Section 3.1). Additionally, we might overlook some commits, although they satisfy our selection criteria although we carefully inspect each collected commits (in Section 3.3).
Threats to external validity refer to the generalizability of our study. In our study, while we involve 1,519 breaking changes extracted from 381 randomly sampled projects from a vast number of popular NPM projects, and these projects cover diverse application fields of JavaScript language, such as utilities, frontend, Web APIs, database, etc., the extracted breaking changes only account for a small proportion of breaking changes in the whole NPM ecosystem. Besides, in this study we do not consider less popular projects since they have short development history and it is less probable to yield some patterns from these projects. However, the BC practices in less popular projects may still differ from those in popular projects. As for another threat, the Reid et al.’s dataset (Reid, 2020) used in our study was released in 2020 (four years ago), and might not include popular JavaScript and TypeScript released after 2020. We manually checked 500 NPM libraries with the largest number of dependents and found that between them [^29] We use the list in https://leodog896.github.io/npm-rank/PACKAGES.html, which is updated automatically (We accessed the page in September 2024). only 16 were initially published after 2020. Hence Reid et al.’s dataset probably covers most popular JavaScript and TypeScript projects, and this threat might not affect the generalizability too much.
Threats to conclusion validity relate to insufficient high-quality BC commits to support our findings. To mitigate the threats, we try our best to extract breaking changes from commit messages, issues, pull requests and changelogs to construct our breaking change dataset. We also carefully inspect our collected breaking changes and retain those related to JavaScript source code changes and explicitly documented by developers, which can help increase the quality of our dataset.
@@ 7.3. Breaking Change Detection and Analysis on Other Platforms
Besides JavaScript, a lot of studies focus on detecting and analyzing breaking changes for other programming languages, especially Java and Python, and their techniques may be learned by breaking change detection for JavaScript. Brito et al. presented APIDiff (Brito et al., 2018) that can detect syntax-related breaking changes in Maven projects, such as method removal and visibility loss by reusing the refactoring detection tool RefDiff (Silva and Valente, 2017). Some open-source tools can also check Java syntactic breaking changes, such as Clirr (Cli, [n. d.]) and RevAPI (Rev, [n. d.]). For Python language, Du et al. proposed AexPy (Du and Ma, 2022) that can detect similar types of breaking changes like module removal and addition of required parameters, which extends the existing tool PyCompat (Zhang et al., 2020) and Pidiff (pid, [n. d.]). Regarding non-syntactic breaking changes, to the best of our knowledge, Zhang et al. proposed Sembid (Zhang et al., 2022) to detect behavioral breaking changes by measuring the semantic difference of call graphs between old and new programs: if a code change’s semantic difference is larger than a threshold and not identified as benign change, it is classified as a BC. However, their used semantic diffs reflect the structural changes of programs, while many semantic breaking changes do not need many of those changes, e.g., only changing one condition, and adding a small change to the output string. Additionally, as the authors have pointed out, Sembid cannot distinguish semantic breaking changes from non-breaking changes such as re-implementation of a method. Therefore, directly adapting Sembid to JavaScript projects is not suitable.
Additionally, with the help of breaking change detection techniques, many works analyzed the impacts of breaking changes to downstream projects. For example, Jayasuriya et al. (Jayasuriya et al., 2023, 2024) investigated the impacts of breaking changes (especially behavioral breaking changes) to client applications in the Maven ecosystem. They concluded with many findings, e.g., 11.58% of the dependency updates contain breaking changes that impact the clients and 2.30% version updates have behavioral breaking changes that impacted client tests. Similar to the findings in Venturini et al.’s study (Venturini et al., 2023), while the breaking changes are only detected in a small number of code changes or version updates, they still cause negative impacts to downstream projects.